OpenStudy (anonymous):

The sum of three consecutive EVEN integers is 42. What are the integers? (Name all three integers.)

OpenStudy (anonymous):

Let the middle number be n. That means\[(n-2)+n+(n+2)=42 \implies 3n=42 \implies n=14\]From that, we discern that the three numbers are 12, 14, and 16.

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

Can you do a similar problem by yourself now? Do you get it, or did you just trick me into doing your homework?

OpenStudy (anonymous):

Three consecutive even integers... x = 1st integer x + 2 = 2nd integer x + 4 = 3rd integer The three integers equal 42 Lets substitute the three integers in ... (x) + (x + 2) + (x + 4) = 42 (now combine like terms...) 3x + 6 = 42 (subtract both sides by 6 to get 3x by itself) 3x + 6 - 6 = 42 - 6 3x = 42 - 6 3x = 36 (divide by 3 to get x by itself) 3x/3 = 36/3 x = 12 Now you can figure out the other two integers by substitution... x = 12 (first integer) x + 2 = 12 + 2 = 14 (second integer) x + 4 = 12 + 4 = 16 ( third integer) check... 12 + 14 + 16 = 42 42 = 42 (correct) ANSWER : The consecutive, even integers are 12, 14, 16

OpenStudy (anonymous):

Very comprehensive explanation kelliegirl33!!

OpenStudy (anonymous):

not a problem...just trying to help